2. Two pulleys 10 in. diameter and 25 in. diameter is connected by an open belt with center distance of 40 in. Find the belt length needed to connect them.

Please answer with detailed solution.


1
Expert's answer
2022-03-22T07:50:03-0400

Belt length for open belt connection


"L = 2C+\frac{\pi(D+d)}{2}+\frac{(D-d)^2}{4C} =2\times 40+\frac{\pi (25+10)}{2}+\frac{(25+10)^2}{4\times 40}=135.75" in